OpenStudy (anonymous):

What is the 9th term of the geometric sequence 4, -20, 100, ...?

OpenStudy (campbell_st):

you need to find the common ratio to do this compare the terms \[\frac{T_{2}}{T_{1}} = \frac{T_{3}}{T_{2}}... or ... \frac{-20}{4} = \frac{100}{-20}.. so .. r = ?\] then the formula for a term in a geometric sequence is \[T_{n} = ar^{n -1}\] n = number of terms ( in your question its 9) a = 1st term ( in your question its 4) r = common ratio which you found above. hope this helps

OpenStudy (anonymous):

thank you